OpenStudy (anonymous):

The surface area of two similar solids are 340yd^2 and 1,158yd^2. The volume of the larger solid is 1,712yd^3. What is the volume of the smaller solid? A) 170yd^3 B) 754yd^3 C) 272yd^3 D) 928yd^3

OpenStudy (anonymous):

x*y=340; (n*x)*(n*y)=1158; n^2=1158/340; n=sqrt(1158/340) (n*x)*(n*y)*(n*z)=1712; n^3*x*y*z=1712; x*y*z=1712/n^3; x*y*z=1712/(sqrt(1158/340))^3=272yd^3